### v2

# Topological Data Analysis in smart manufacturing

Abstract
Topological Data Analysis (TDA) is a discipline that applies algebraic topology techniques to analyze complex, multi-dimensional data. Although it is a relatively new field, TDA has been widely and successfully applied across various domains, such as medicine, materials science, and biology. This survey provides an overview of the state of the art of TDA within a dynamic and promising application area: industrial manufacturing and production, particularly within the Industry 4.0 context. We have conducted a rigorous and reproducible literature search focusing on TDA applications in industrial production and manufacturing settings. The identified works are categorized based on their application areas within the manufacturing process and the types of input data. We highlight the principal advantages of TDA tools in this context, address the challenges encountered and the future potential of the field. Furthermore, we identify TDA methods that are currently underexploited in specific industrial areas and discuss how their application could be beneficial, with the aim of stimulating further research in this field. This work seeks to bridge the theoretical advancements in TDA with the practical needs of industrial production. Our goal is to serve as a guide for practitioners and researchers applying TDA in industrial production and manufacturing systems. We advocate for the untapped potential of TDA in this domain and encourage continued exploration and research.
keywords:
Industry 4.0 , Manufacturing , Topological Data Analysis , Persistent Homology , UMAP , Mapper
journal: a Journal
[…]
1 Introduction
Industry 4.0 represents the fourth industrial revolution, marked by the integration of digital and physical technologies. It is reshaping manufacturing and facilitating the creation of smart production systems. Such systems are distinguished by their geographic distribution, interconnectedness, and capability for autonomous decision-making to swiftly adapt to changes in production requirements [1]. Traditionally, these adaptations were managed by operators, a process requiring extensive training and multiple attempts to produce acceptable goods.
Topology, a branch of mathematics, focuses on studying the properties of objects that remain unchanged under continuous deformations. Topological Data Analysis (TDA), an emerging field at the confluence of data analysis, computer science, and algebraic topology, leverages the underlying topological and geometrical structure of datasets to uncover features not detectable with other analysis methods. This approach enables the analysis and visualization of topological structures such as connected components, loops, and cavities, and has been proved efficient across various applications including anomaly detection [2], image processing [3], and genome sequencing [4], as well as in sectors like the chemical industry [5], finance [6], aviation [7], and physics [8].
The shared challenges and characteristics across different fields highlight the substantial promise of TDA in industrial production contexts. Yet, the integration of TDA within industrial manufacturing is notably limited. We claim that TDA is particularly effective for intelligent production systems, as it facilitates the extraction of valuable insights from complex data generated by sensors and various devices, thus supporting enhanced decision-making. However, its comprehensive potential in the sphere of Industry 4.0 has not been fully realized.
1.1 Contribution & Outline
This paper offers a comprehensive review of the current literature on the applications of TDA in industrial production and manufacturing. It aims to bridge the existing gap in mutual awareness between theorists in TDA and practitioners in industrial production, facilitating the exchange of ideas and methods between these domains. Since the theoretical nature of TDA may be off-putting, we provide a hands-on description of its tools and exemplify them with industrial applications.
This paper makes the following four contributions:
  1. 1.
    It provides an overview of the current literature on TDA applications within industrial production and manufacturing;
  2. 2.
    It demonstrates how TDA methods are currently being integrated into smart manufacturing;
  3. 3.
    It highlights areas within smart manufacturing where TDA methods are underexplored;
  4. 4.
    It outlines some guidelines for off-the-shelf uses of TDA in industrial production and manufacturing.
The structure of this paper is as follows: Section 2 introduces the necessary background and key terminology related to Industry 4.0. Section 3 presents TDA and its principal tools. The methodology of the survey is outlined in Section 4, followed by the presentation of results in Section 5, which includes a comprehensive discussion for each identified application domain. Section 6 interprets the findings and suggests directions for future research. The paper concludes in Section 7.
1.2 Related Work
The application scope of TDA has expanded so significantly, encompassing over 450 papers to date, that it has become virtually impossible for a single researcher to keep up with all the advancements. This proliferation raises the question of how to structure the interface between TDA and its application domains to make it accessible to both theorists and practitioners.
The platform DONUT, introduced by Giunti et al. [9], offers a search engine designed to facilitate the exploration of TDA applications. Additionally, scientific surveys play a crucial role in organizing this knowledge, summarizing and contrasting different approaches applying TDA in practical settings. There exists a considerable collection of such surveys (e.g. [10, 11, 12, 13, 14, 15, 16]) and textbooks (e.g. [17, 18, 19]) that generally focus on explaining the theory and showcasing exemplary applications. Nevertheless, we found none focusing on industrial manufacturing. There are, however, reviews on TDA applications in other fields, such as Topological Time Series [20] or Topological Machine Learning (ML) [21, 22, 23], as well as reviews on applications in manufacturing without a TDA focus.
Capodieci et al. [24] review ML and other data analysis methods for yield optimization in semiconductor manufacturing, briefly mentioning Mapper’s application. Gao et al. [25] discuss the potential of TDA methods, including Mapper, for Big Data analytics in smart factories. Kounta et al. [26] and Liewald et al. [27] highlight TDA’s potential in machining chatter detection and metal forming, respectively.
Additionally, Big Data’s relevance to Industry 4.0 is underscored in a survey by Snasel et al. [28], aiming to bridge theoretical and engineering disciplines. Wang [29] discusses Persistent Homology (PH) and Mapper in 3D printing, a key technology in additive manufacturing.
2 Smart Manufacturing in Industry 4.0
2.1 Industry 4.0
The term Industry 4.0 was coined in the early 2010’s by the German government in order to foster and drive the so-called “fourth industrial revolution” (see [30] for details). This movement is driven by the need for higher flexibility in production, the operation of machines that are more adaptive to the requirements of the production, and a smarter, more autonomous way of operating machines, production lines, factories and even whole supply chains. This enables production paradigms like lot-size-one production and mass customization. However, it also fosters the optimization of production scenarios not only per machine but on the entire value chain, eventually enabling new business and operational models as well. This vision gives rise to various other terms along with Industry 4.0, like “smart factory” or “cognitive factory”.
However, Industry 4.0 is not only about manufacturing of products only, but its context is much wider. Indeed, [30] argue that this paradigm enables more innovative and revolutionary products per se by incorporating other technologies that are occurring simultaneously. These technologies include gene sequencing, nanotechnologies, renewables, quantum computing, and much more [30].
Hermann et al. [31] identified the following four design principles of Industry 4.0 systems:
  • 1.
    Interconnection: all components, like sensors, machines, and even humans, are connected with each other.
  • 2.
    Information Transparency: the information about all components are transparent within the system. This enables operators to make intelligent and well-informed decisions.
  • 3.
    Technical Assistance: the technological facilities assist humans in decision-making, support in problem-solving, and help or take over hard or unsafe tasks.
  • 4.
    Decentralized Decisions: decisions are not made by a central instance, rather they are made “on the edge”. Cyber-physical systems are able to make decisions on their own, based on the information they have. Exceptions to this rule are in inferences or conflicting aims, delegated to a higher instance.
A review by Erboz [32] identified main components of systems of Industry 4.0: Big Data and Analytics, Autonomous Robots, Simulation, Horizontal and Vertical System Integration, Industrial Internet of Things (IIoT), Cloud, Cybersecurity, Additive Manufacturing, and Augmented Reality (AR). When talking about Cybersecurity in the context of Industry 4.0, or Operational Technology (OT) in general, we further refer to the term OT security [33]. The term Additive Manufacturing refers to the technology of 3D Printing in an industrial production context. Here, three-dimensional objects are created layer-wise by deposing material in a computer-controlled process [34].
2.2 Production and manufacturing
In the literature, the terms Manufacturing and Production are used for the manufacturing process of producing goods. Depending on the domain also other terms are found, e.g., Fabrication in semiconductors [35]. Despite slight semantic differences, we use the term Production and Manufacturing as umbrella terms for these three terms within this survey.
The manufacturing of a product involves a sequence of process steps, applied by industrial machines in a production line. Figure 1 illustrates the stages of the manufacturing engineering process sequentially. At the beginning of a design process, the definition of the product requirements have to be identified, followed by a conceptional design and evaluation of the said. Based on that, a prototype is created, enabling the creation of schematics for industrial reproduction, where “industrial” typically means in a repeatable, efficient, and effective way. These schematics, in combination with the requirements of the product, define the specification for the selection of materials, processes, and production equipment. The production itself is then accompanied and finished by an inspection and quality assurance before the products are packed.
Manufacturing- or Production Engineering describes the branch of engineering working on the entire process of manufacturing. Among others, the planning and optimization of the production process are subject of interest to this discipline [36].
Figure 1: The manufacturing engineering process with its stages along product production, beginning from the productions definition to its final, mass-produced artifact. Feedback connections omitted for the sake of readability. Illustration adapted from [35].
In general, smart manufacturing in Industry 4.0 poses new challenges to the domain, compared to traditional manufacturing. Here, additional strategies and technologies are used to improve the manufacturing process, to fulfill the needs for the integration into Industry 4.0. An overview on the technologies and architectures used for smart manufacturing systems is given in [37].
3 Topological Data Analysis
We consider TDA to consist of three main methods: Mapper [38], PH [17, 39], and Uniform Manifold Approximation and Projection (UMAP) [40]. PH is the prevalent method in TDA, Mapper a topologically guided clustering algorithm, and UMAP a topology-driven dimensionality reduction method. Common to all of them is that (i) the data at hand is converted into a suitable (geo)metric representation and (ii) topological properties are analyzed to extract (novel) information from the given representation. In a remote analogy, where the Fourier analysis tells us something about spectral properties of a signal, TDA extracts information about the shape of various kind of data (e.g., loops, tunnels, cavities, components, et cetera).
The high-level pipeline of the three methods is depicted in Figure 2: input data is transformed into a combinatorial object, which is then used to extract and interpret information. This is done either directly (Mapper and PH) or by feeding its (vectorized) output to a ML pipeline (all three methods).
Figure 2: General TDA pipeline. Images of the time series (real and simulated) are taken from [41, Figures 4 and 9], respectively. The simplicial complexes and the point cloud images come from [42, Figures 1 and 3], respectively. The graph and the clustered graphs are taken from [38, Figure 5] and [40, Figure 2], respectively.
During the stage of data preprocessing, data usually is manipulated by a set of parameters. The three aforementioned methods deal with parameters – i.e., the step from input data to a combinatorial object – in different ways. First, Mapper assembles parameters (and their values) in different groups and then clusters the inputs accordingly. This grouping can find latent relations in the dataset. Similarly, UMAP is a dimensionality reduction method. By projecting the data into lower-dimensional ambient space, redundant or not relevant information in the data is compressed. This enables the input to be analyzed more easily. On the other hand, PH overcomes the need of choosing parameter values by considering all possible values at the same time. Instead of outputting information about the shapes in the data for a single parameter value, PH keeps track of how the shapes evolve depending on the variation of parameter(s). PH is especially suitable for automated production since it removes the need to cherrypick a threshold (see for example [43]).
A considerable attribute common to all TDA methods is that they are applicable to virtually any scenario: as long as the data can be converted to some general formats, the methods are agnostic to the data’s source. Therefore, TDA can be applied to any domain that has data in a suitable input form or that can be transformed into it, such as point clouds, meshes, polygonal shapes, time series, graphs, pictures, and more. Domain-specific knowledge is only needed at a later stage to interpret the results.
Figure 3: Scheme of the described TDA methods, with pro, contra, and (dis)similarities.
In Figure 3, we summarize the advantages and disadvantages of the three aforementioned methods and highlight the dissimilarities and shared characteristics.
3.1 Mapper
Mapper [38] reduces the data dimension by clustering the input points in a graph, the so-called Mapper Graph. The graph’s vertices represent the clusters and the edges signify that two clusters share some elements. This representation can further be analyzed for its topological or geometrical properties, such as the connected components, loops, and flares.
Figure 4: Mapper illustrated differently, indicating the choices the user has to make.
Figure 4 outlines a high-level overview of the Mapper pipeline, highlighting the necessary choices to be made. In detail, three choices are required:
  1. 1.
    a real-value function $f\colon X\to\mathbb{R}$ over the input data, called filter or lens function,
  2. 2.
    an interval cover, cover for brevity, of the image of $f$ (i.e., a collection of possibly overlapping intervals whose union covers the image of $f$), and
  3. 3.
    a clustering algorithm.
These choices are highly application-dependent and different choices usually result in different outputs (and therefore in different interpretations).
The filter function may reflect geometrical properties of the data (for example, a density estimator) or some wanted feature the user is interested in (for example electrical properties). In theory, Mapper works for general filter functions, e.g., maps to the circle $S^{1}$ or $\mathbb{R}^{d}$. In practical applications, the filter function is typically a real-valued function due to its simplicity and proven efficiency.
Figure 5: Graphical illustration of the Mapper Algorithm.
We now describe the steps of Mapper (illustrated in Figure 5): first, the input data $X$ is reduced in dimension using the filter function $f$. Then one chooses a cover of the image of $f$. Once the image of $f$ is covered, one takes the preimage of each interval in the cover, i.e., the collection of all data $x\in X$ such that $f(x)\in\mathcal{I}$ for an interval $\mathcal{I}$ in the cover. In other terms, all data points that are mapped to the same interval in the cover are grouped together in the domain. Note that, since the intervals in the cover may overlap (and they usually do), a point $x\in X$ can belong to the preimages of several intervals. Each of these preimages is clustered according to the chosen clustering algorithm. Finally, each of these clusters is represented as a vertex, and two vertices have an edge between them if there is at least one $x\in X$ that belongs both to corresponding clusters.
The Mapper graph is used for explorative data analysis: for example, one looks for flares in the graph, that is, subpopulations of vertices that are connected to each other across several intervals and disconnected from all others vertices. One then analyzes these subpopulations (potentially with traditional data analysis methods) to find a reason for their distinctiveness. For some practical examples outside industry, see [44]. Moreover, the Mapper graph can be feed to a ML pipeline. Implementations of Mapper are freely available (see, for example, [45, 46]).
To give an idea of how the method works in practice, we now describe a practical example taken from industrial manufacturing (c.f. Section 5). The authors of [47] develop a demand-forecasting system of electronic components that uses Mapper to identify which forecasting model better predict the demand of a given component. This system is constructed in two steps: first a Mapper Graph is built using a training subset of components, and then the most fitting model for a new component is selected using this graph.
The input data are time series encoding the demand of a given electronic component over a certain time span. A small fraction of the data is used as the training set. This fraction is preprocessed by computing the features describing each time series and its best forecasting model. The collection of time series features is the actual input for Mapper, which maps them using the first principal component as a filter function. After covering the image and clustering the preimages, we obtain a graph whose clusters represent electronic components whose demands behave similarly. These clusters are labelled with the forecasting models that works best for (the majority of) the electronic components in it. The resulting Mapper graph can then be used to predict the best model for new components.
This prediction of a new, unknown component is performed as follows: the components features are calculated (as for the training sets). Then a $k$-nearest neighbor search is performed on the Mapper graph to find the nodes with similar features. The cluster with the lowest distance to the new component is assigned to it. The forecasting model of the assigned cluster is then applied to the new component.
In practice, the major obstacles to the use of Mapper are the choices of the filter function $f$, cover, and clustering algorithm: the interpretability of the outcome entirely depends on these choices. While a few standard choices are known, a domain expert is usually required to get meaningful insights out of the Mapper pipeline. Nevertheless, Mapper is a powerful, versatile method that can reveal covert connectivity in datasets.
3.2 Uniform Manifold Approximation and Projection (UMAP)
UMAP [40], as Mapper, is a topologically driven dimensionality reduction method. It works by building two (weighted) graphs. The first graph is embedded in the same ambient dimension as the data and approximates the geodesic distances of the underlying manifold. The second graph approximates the first graph in a much lower ambient dimension. The data’s embedding in this lower dimension is then used for further analysis. Figure 6 illustrates UMAP schematically.
Figure 6: UMAP pipeline. High/low-dimension refers to the ambient space where the graph lives.
The first graph is constructed in several steps: First, the $k$-nearest neighbors of each point $x_{i}$ in the input data $X$ are computed. Here, $k$ is a parameter chosen by the user (different values of $k$ may result in different outcomes). Then the geodesic distance from $x_{i}$ to each of its neighbor is approximated. All distances for each input point $x_{i}$ are aggregated. These distances provide an incomplete and conflicting collection. Indeed, take two points $x_{i},x_{j}\in X$. The approximated distance from $x_{i}$ to $x_{j}$, computed while processing $x_{i}$, is not necessarily the same as the one from $x_{j}$ to $x_{i}$, computed while processing $x_{j}$. Moreover, while studying the $k$-nearest neighbors of $x_{i}$, we get the approximated distance from it to $x_{j}$ and $x_{h}$, yet we have no information about the distance between $x_{j}$ and $x_{h}$. To overcome this issue, the authors use fuzzy structures and directed graphs. Fuzzy structures are a way to weight the membership to a set (instead of either being in or out, an element can belong to a set fuzzily). The membership strength is given, in this case, by the approximated distances. The directionality in the graph takes care of the asymmetry of the distances. We omit further technical details, which are not relevant for this survey (see [40] for details).
Since the distances are approximated from each point independently, the computation can be parallelized and it scales well with large inputs. Another advantage of this construction is its robustness against the curse of dimensionality, given by the fact that the distances are approximate locally.
The second graph is constructed by projecting the first graph into a lower dimensional manifold. This projection is achieved via a force directed graph layout algorithm, which requires the user to choose two further hyperparameters. Contrary to what happen with other methods of dimensionality reduction, such as t-distributed stochastic neighbor embedding (tSNE) [48], an arbitrary target dimensionality can be selected. The resulting second graph encodes the original local distances from the input, but lives in a lower dimensional ambient space. This final representation can be more easily processed with other data analysis methods, in particular ML.
There are several free implementations available of UMAP, such as [49, 50, 51].
The work by Hsu et al. [52] exemplifies the application of UMAP. They analyze failures in System Level Tests (SLTs) on integrated circuits. The need for robust interpretative methodologies is driven by the increasing transistor density and complexity in circuit designs, where the high-dimensional nature of the associated data poses a significant challenge. To address this, Hsu et al. [52] project high-dimensional data into a lower-dimensional space using UMAP. The obtained lower-dimensional representation is analyzed with a clustering algorithm, isolating clusters of systematically failing SLTs. These clusters are then leveraged to construct a decision tree, enabling precise identification of failures in individual components. The clustering algorithm is very efficient but sentitive to the curse of dimensionality. Therefore, applying it directly to the data would lower the performances.
Contrary to the outputs of Mapper and PH, the output of UMAP is usually not immediately interpretable, but can be used for further analysis, e.g., using ML. In conclusion, especially since it scales well with bigger input, and its robustness against the curse of dimensionality, UMAP is a powerful tool for dimensionality reduction that has proven to be very successful in practice, outside and inside industry.
3.3 Persistent Homology (PH)
Many data preprocessing techniques require a fixed choice for the value of one or more parameters. The key idea of PH is to use topological methods to study how features in data evolve for different choices of the parameter. For instance, rather than choosing a fixed threshold for binary thresholding in image analysis, we consider the sequence of results over a sweep of thresholds. Or instead of considering clusters of a point cloud by investigating the union of balls of fixed radii centered at the points, we consider the evolution of this union over a sweep of radii. This not only allows us to avoid the choice of a parameter, but also to extract more information from the data, by adding a dimension to the analysis.
In the case of PH, the combinatorial objects representing the mentioned evolution of data are simplicial complexes. These can be seen as a higher-dimensional generalization of graphs. While graphs only have edges that link two vertices, simplicial complexes can link $n$ vertices and thus form edges, triangles, tetrahedrons, and so on (for some examples, see Figure 7).
Figure 7: Three simplicial complexes, each including into the following forming a filtration.
Contrary to what happens for Mapper and UMAP, where most of the work goes into building the combinatorial object, for PH this is just half of the effort (see Figure 8). To get to the combinatorial object from the input data $X$, one first has to choose the varying parameter $r$ [^1] For the sake of simplicity, we restrict to the case of just one parameter; there is a rich theory for multi-parameter persistence, but the pipeline becomes more complicated and many successful applications work with just one parameter, where we also have more efficient algorithms.. Then, for each value of $r$, one obtains a space $X_{r}$ that results from the input data using $r$ as the parameter value and turning it into a simplicial complex (we can think of it analogous to a triangulation).
PH considers the evolution of $X_{r}$ with increasing parameter $r$. Here we require that $X_{r}\subseteq X_{s}$, when $r\leq s$. The sequence of $X_{r}$ is called a filtration, and it is the combinatorial object investigated by PH.
There are several methods to turn $X_{r}$ into a simplicial complex. We will not elaborate on details, but this choice may affect the interpretation and usually impacts the runtime considerably. Therefore, it is important to choose the more appropriate method for each application to ensure efficient computation [53].
Figure 8: PH pipeline.
Now that we have a suitable combinatorial object, we need to extract information from it and we do so using homology. Homology is a fundamental concept from algebraic topology, allowing us to identify shapes that cannot be continuously deformed into each other. An extensive treatment is beyond our scope (see for example [17] for more details). Informally, homology reveals the number of $k$-dimensional holes of a shape, for every integer $k$. For $k=0,1,2$, this corresponds to the number of connected components, tunnels, and cavities in the shape. Crucially, given a continuous map between two shapes (or, in this case, simplicial complexes), for instance, an inclusion from $X$ into $Y$, there is a well-defined map between the holes. Therefore, we can study the evolution of the homology in the filtration we obtained from data. The resulting evolution of topological features can be represented as a barcode (aka persistence diagram), a collection of intervals (bars) that represent the lifetime of a hole within the filtration. The length of a bar is called the persistence of the corresponding topological feature.
For a practical example, we describe the application presented in [42]. Fine-grained segmentation of 3D images is the labeling of small components of objects. This is crucial in object analysis such as intelligent manufacturing, autonomous robotic manipulation, and general automatisation procedures. To obtain correct fine-grained segmentation, it is important to identify geometric and topological properties at different scales (e.g. shelves, handles, and hinges in a closet). Moreover, often these properties involve more complex structures than the ones obtained by mutual distances. The authors of [42] use PH to enhance the prediction of a convolution network. The input is a 3D shape in point cloud format, and it is processed in parallel by a graph convolution network and PH, which both output a representation in a latent space. These vectors are then combined and augmented to obtain the augmented feature map which label the input data.
PH produces the persistence diagrams of the loops and cavities (1- and 2-dimensional homology, respectively). These diagrams are vectorized using Persistence Images (PIs) [54]. Then, to train the network, the authors [42] use a combination of the (commonly used) cross-entropy function and a newly defined PH loss function. This latter ensures a small topological error, ensuring a much more precise segmentation at different scales.
As illustrated in Figure 8 and exemplified in [42], barcodes can be vectorized and thus fed to a ML pipeline for enhanced analysis. There is a rich theory on how to compare two datasets by comparing their barcodes, and how to integrate PH into ML methods, i.e., kernel-based methods [55, 56] or neural nets [54, 57]. The well-founded theory and the interpretability of the obtained features have contributed to the success of PH in practice. There are many efficient algorithms to compute filtrations and barcodes, and to subsequently compare them, such as [58, 59, 60, 61, 46].
A drawback of PH is that, unlike Mapper or UMAP, it does not easily scale to very big datasets. On the positive side, it is more intuitive than Mapper and UMAP because usually there is a natural choice of parameter to filter. Moreover, since it requires less parameter tuning, PH is very apt for automation.
4 Survey Method
One of the objectives of this work is to ensure the reproducibility of the survey results. Therefore, we decided to perform the review as an exhaustive literature review [62], where each step is documented and can be reproduced. The problem defined for this survey is the review of methods from TDA in industrial production and manufacturing.
The pipeline of the survey method is illustrated in Figure 9 and contains the following steps: definition of appropriate keywords for the search, identification of the digital libraries where to search, and filtering of the obtained works. We now explain these steps in detail.
Figure 9: The data aggregation and filtering pipeline for the literature review consists of several steps. The first step is the definition of a set of keywords for the search (Section 4.1). Further, these keywords are used to search in the digital libraries (Section 4.2). The results are then filtered according to the criterias described in Section 4.3. The blue boxes indicated automated steps, while the green ones indicate manual effort.
4.1 Keywords and queries
The search queries are defined by means of two categories, namely Method and Domain. The keywords of the categories Method describe the TDA tools whose applications we are interested in finding. On the other hand, the keywords of the Domain category describe applications and tasks in industrial manufacturing. Figure 10 illustrates these categories including the identified keywords. The intersection of both categories is the search space for our literature review.
Figure 10: Venn diagram of the set of 13 keywords for the category Domain (left segment) and the set of 4 keywords for the category Method (right). The intersection of both sets indicates the scope for this literature search. Note: the asteristk “*” indicates a wildcard, so that variants of the keyword can be considered (e.g. Technologies and Technology).
The search queries are formulated depending on the search functionality of the digital libraries. Using an individual search query each, the results are collected and stored.
4.2 Digital libraries
We surveyed the following five digital libraries:
  • 1.
    IEEE with the IEEE Xplore Digital Library
  • 2.
    Springer with SpringerLink
  • 3.
    Elsevier with ScienceDirect
  • 4.
    ACM with the ACM Digital Library
  • 5.
    The American Society of Mechanical Engineers (ASME) with the ASME Digital Collection
We chose these digital libraries because the first four are the most prominent digital libraries for scientific publications in the field of computer science and engineering, especially for the applications of TDA and ML. Furthermore, these four are also the most prominent digital libraries for scientific publications in the field of industrial engineering. The only exception from this selection criteria here is the ASME Digital Collection. It was selected because an earlier and preliminary semi-exhaustive search – using a method similar to the one used in [63] – on a restricted set of keywords on Google Scholar showed additional relevant results that are contained in the ASME Digital Collection but not in the first four collections.
IEEE Xplore, SpringerLink, and the ACM Digital Library provide a search interface that allows the usage of the generated search strings. The interfaces of ScienceDirect and the ASME Digital Collection limit the length of search queries, so we split our search query into several sub-queries and aggregate the results. From all these digital libraries, all results are collected and stored using the Zotero reference manager [^2] https://www.zotero.org/. For each publication its metadata is automatically extracted through its Digital Object Identifier (DOI).
4.3 Filtering results
For the filtering, we did not restrict the search results on a specific time period. The reason for this is that TDA is still a young field and the application of it in industrial manufacturing is even younger. The data aggregation was performed on April 23rd, 2024, and all published works up to that date are considered.
The filtering procedure was done step-wise, removing items according to the following criteria:
  1. 1.
    Duplicates
  2. 2.
    Relevant types of publication
  3. 3.
    Language of the references
  4. 4.
    Availability of a full text
  5. 5.
    Context of the work
The first step of the filtering procedure was to remove duplicates, where duplicates are identified by the DOI and the title of the publication. When duplicates were found based on the title then a manual check was performed to verify this duplicity. Each duplicate was removed, keeping only one instance of the publication [^3] Here to be noted, Conference Paper with a following, extended Journal version are considered to be independent (since they have a uniqe DOI, respectively). Their relationship is indicated in the results section (Section 5)..
In order to deliver a meaningful review, only publications with a sufficient degree of quality were considered (i.e., peer-reviewed publications). Based on this requirement, only publications from conference proceedings and journals were taken into account for our survey. While not all conference and journal publications are peer-reviewed, the majority of them are, and hardly no status on a review process is actually provided by the digital libraries. Other references, like preprints, presentations, books, or reports are excluded and not further considered for our analysis.
To ensure the correct extraction of information from the publications and to be commonly reproducible, only publications with an available English full text were considered. The availability is highly dependent on the access and subscriptions of our institutions to those digital libraries. Therefore, all publications with no available full text were screened manually, so that no relevant publications are missed.
The semi-automatically filtered references are then analyzed for the context. For this, all publications were screened manually. Here, the keywords of the categories Method and Domain were searched in the publications. They had to be present in a section relevant to the contribution. In particular, it was not sufficient to mention any of the methods or domains only in the related work or in the outlook.
After this filtering procedure, only the remaining publications are further considered for this survey. In total, over $7000$ results were screened, resulting in $35$ publications that were considered to be relevant for our stated research questions. These publications were then screened in detail. Based on this screening, the publications were categorized manually into different categories, that are presented in the following.
5 Results
In total $35$ works were identified as relevant for this survey. Each of those identified works was assigned to one of three clusters (A-C), based on its application within the production process. The identified clusters are:
  • 1.
    A: Quality Control on Product Level
  • 2.
    B: Quality Control on Process Level
  • 3.
    C: Manufacturing Engineering
The resulting works are listed in Table 1. This table indicates each work with its assigned cluster (A-C) and the used TDA method.
Cluster TDA Methods PH Mapper UMAP A  [64, 65, 66, 67, 42, 68, 69, 70, 71]  [72]  [73, 74, 75] B  [76, 77, 78, 41, 79, 80, 81]  [82, 83] C  [84, 85, 86, 87]  [47]  [52, 88, 89, 90, 91, 92, 93, 94]
Table 1: The results of the survey with all references, assigned to a cluster (A-C) over the applied TDA methods (PH, Mapper, and UMAP). A more detailed representation is shown in Table 2.
Figure 11 provides an overview of these works in relation to their assigned clusters. Furthermore, this illustration also indicates the used TDA method over each application area.
Figure 11: This illustration shows the relation of the works to the identified clusters. Furhtermore, the used TDA method is indicated for each cluster. The number in the parenthesis indicates the number of publications associated.
From the publication dates of the listed works can be seen that the interest in TDA methods for manufacturing has increased over the last years. The first articles on the topic were published in 2016. A major gain in interest can be observed for the years 2022 onwards. The count of publications in 2022 is higher compared to the number of publications in the years before 2022. Similarly, the number of publications in 2023. Figure 12 shows the cumulative count of relevant publications per year. Note that the data for 2024 is incomplete, as this survey includes publications added to the digital libraries up to April 23rd, 2024. From this plot, we can see that the interest on TDA methods in manufacturing has increased in the last years. Similar trends can be observed in other fields of application, see for example [95, 96].
Figure 12: The number of relevant publications per year. Counts of all methods for each year are stacked, displaying a cumulative total. Data aggregation was performed on April 23rd, 2024. Information regarding the publications is available in Table 2.
A more detailed overview of the results is given in Table 2. This table shows each individual work, the associated cluster, the used TDA method, and the kind of input data used to solve the task. The type of input data is extracted from the referenced works. The most common data type is time series data, followed by point clouds and scalar fields. Additionally, we found also one work employing the TDA method on textual log files and one on a labeled graph.
Cluster Work Input Data Format fin. $\subset\mathbb{R}^{n}$ $\mathbb{N}\to\mathbb{R}^{n}$ Scalar Field ($\mathbb{R}^{2}\to\mathbb{R}$) Graph Point Cloud Time Series Wafer Map Sign. Dist. Func. Microscope Imag. Surface Logs Funct. Block Netw. Task Graph A  [42] PH  [64] PH  [66] PH  [67] PH  [65] PH  [68] PH  [69] PH  [70] PH  [71] PH  [72] Mapper  [73] UMAP  [74] UMAP  [75] UMAP B  [41] PH  [76] PH  [77] PH  [78] PH  [79] PH  [82] Mapper  [83] Mapper  [80] PH  [81] PH C  [47] Mapper  [52] UMAP  [84] PH  [85] PH  [86] PH  [88] UMAP  [89] UMAP  [90] UMAP  [91] UMAP  [92] UMAP  [87] PH  [93] UMAP  [94] UMAP Count $7$ $17$ $3$ $1$ $2$ $2$ $1$ $1$ $1$ $8$ $2$
Table 2: All the identified works of the survey including their data format employed on and the method being used, grouped by the identified clusters. The data format is expressed in a mathematical notation and their occuring format within the works. The works on time series data are of different kind and employ the TDA methods to uni- and multivariate time series data. As scalar fields are considered Wafer Maps (also Wafer Defect Maps), Signed Distance Functions, Microscope Images, and Surfaces. Logs are considered to be heterogeneous and multidimensional textual data, with the necessity of preprocessing this unstructured format. The functional block network is a labeled graph, where the nodes are functional blocks and the edges are the connections between the blocks. The task graph is a graph, where each node represents a task and the edges are weighted according to the connected tasks properties.
In the following Sections 5.1, 5.2 and 5.3, the identified application clusters are discussed. For each area of application, a short description is presented, followed by a brief summary of the associated works. For further details on these works, we refer to the original, referenced publications. Additionally, Section 5.4 discusses works encountered during the survey that apply TDA methods to validate analyses performed by other methods.
5.1 Quality Control on Product Level
This cluster encompasses works focused on identifying discrepancies in manufactured goods. Different tasks were identified, but all share the common goal of comparing a measured property with a reference. During the production process, this task is performed after the production itself, as illustrated in Figure 1 (cf. Inspection and quality assurance). In general, methods from TDA are naturally suited for the analysis of structures, surfaces, and shapes. We survey the usage of TDA in the context of quality control on product level.
A natural application of TDA on the product level is the analysis of discrepancies of the products topology, as shown in the example in [64]. In this work, the authors classify topological discrepancies in additive manufacturing using TDA methods. The products to be classified are embedded as a mesh in $\mathbb{R}^{3}$. In this work, they mainly use (non-persistent) homology [^4] For the sake of simplicity, we account this work as an application of PH, as mathematically PH can be considered to be a generalization of homology.. Discrepancies are detected by comparing the homological features of the mesh with the homological features of a reference mesh. The authors are able to demonstrate the successful application of their method to the classification of deviations between as-design and as-produced articles.
Another natural use case is the analysis of surfaces texture. An early work proposes the application of PH to differentiate various generic surface profiles [66]. Their method applies the filtration on the captured profiles, describing the surface lattice by 00- and $1$-dimensional persistent homology features. In a follow-up work [67], the authors apply this methodology on the more specific task of microscope images, recorded after a Piezo Vibration Striking Treatment. This treatment is a method to change the surface finish of a product, which is very essential for e.g. its hardness. With both works, the authors are able to report similar results as their baseline methods, while eliminating the manual preprocessing steps, demanded by traditional methods. In [65], the authors extended their approach to further analyze additional properties. The “tool striking depths” and the “roundness” of the patterns onto the more general format of surface images are analyzed in addition to characterizing the patterns on the product’s surface.
For the segmentation of shapes in point cloud data, [42] propose a new method employing PH and graph convolutional networks. Their approach enriches the “local” features, extracted using the graph convolutional network, by “global” features. These features are extracted using PH, transformed to PIs and reshaped to a “topological feature map”. Their PH-based graph convolution network outperforms the state of the art on methods of fine-grained 3D shape segmentation, performed on point cloud data.
A more specialized use case is the identification of defect patterns in the wafer production. For this task, [72] case-study the use of Mapper. Their method takes wafer maps as the input. In a first step, a vision transformer extracts higher-level features from the wafer map images. Mapper is then applied on these extracted features. The Mapper graph’s visualization is then used to identify clusters. the authors highlight their method’s ability to identify defect patterns on a public benchmark dataset, without a comparison to other methods.
Another application scenario is the production of electrical motors. In [68], the authors use PH for the detection of eccentricity of electric motors. Eccentricity is a non-uniform gap between the stator bore and the rotor that causes a non-circular rotation. In this task, the input data is a time series of the process parameters of the electric motors. The time series is transformed into a point cloud in 3D, using a so-called time-delay embedding. On this transformed data representation, PH is applied. The eccentricity level can then be described by the 1-dimensional homological features in the persistence diagram. In their work, they are able to predict the fault levels with a reasonable accuracy, while achieving significantly lower computational complexity. In the follow-up work [69], the authors compare their method to physics-based approach, and in [70] the authors provide a more thorough theoretical description and experimental evaluation of their method.
Similarly, [73] analyzes the root cause in manufacturing variations of mechanical components during a production process. For this task, the input data is a 3D point cloud, acquired from an optical scanning step. With their work, they apply UMAP as a manifold learning method towards this task.
The work in [74] proposes a method for the sorting of electrical wafers. The primary objective for their use case is to detect defects in produced wafer maps, so that these can be discarded at an early stage. Within the proposed Deep Learning pipeline a preprocessing step is performed. This preprocessing applies a dimensionality reduction on the high-dimensional input data. For this task, they employ UMAP as a dimensionality reduction method.
A similar task is performed in a work [71], where the authors detect defect patterns in wafers maps. In this work, the authors propose a method using PH. First, a filtration is performed on the wafer maps in $\mathbb{R}^{2}$. The resulting persistence diagram is then transformed into a vector representation, using PIs. The neural network then performs the classification task on the vectorized representation. Moreover, the authors highlight the computational efficiency of the proposed technique compared to other methods.
On the task of additive manufacturing of radio-frequency devices, [75] propose a method for the identification of defect mechanism and their performance impact. This task usually requires in-line electromagnetic simulations, which are time-consuming and expensive. In their work, they use convolutional neural networks on microscope images, resulting in a vector. UMAP is then employed for the dimension reduction on the resulting vectors. Their approach provides to a faster and cheaper quality control process.
5.2 Quality Control on Process Level
The applications in this cluster aim to assess production process quality by observing the state of the process. These observables are assesed during the production process (cf. Figure 1, Production). Examples of data in this context are machine states, sensor data, or other data acquired during the production process. From these process variables, failures, anomalies, and other quality issues can be detected.
Nine works belonging to this cluster were found during the review. These works deal with two major application types. As illustrated with the example of an injection molding machine in Figure 13, all applications (with one exception) are based on time-series data as the input.
Figure 13: Applications of quality control at the process level utilize key process parameters, represented as time-series data, to monitor and control machinery processes. Process parameters are illustrated here on a schematic injection modeling machine. This figure is complete regarding the input data types (time-series) on the task of this cluster, as shown in Table 2.
Given observations from key process parameters, the aim of the works in [82] and [83] is to predict the productivity of a manufacturing process. The authors use Mapper to identify intrinsic clusters within benchmark processing datasets. Using Mappers output network, key process variables or features that impact the final product quality the most are selected. Their work showed that this model achieves the same level of prediction accuracy as with all process variables, while being more cost-effective.
The second application within the cluster of quality control on process level is the detection of chatter. Chatter detection in machining has garnered attention in recent years, as can be seen by a survey work on this particular application domain [26]. Detection of chatter is important since the abnormally of chattering can lead to damage to the workpiece or the machine tools. Using TDA methods for chatter detection has been studied in a line of research by Khasawneh et al.
The first work [76] is a proof of concept, where the authors show that PH can be used for chatter detection in general. Each time series is converted into a delay embedding. The structure within this representation can then be extracted using PH. Abnormalities are then visible within the persistence diagram.
In the follow-up work [77], the authors propose a method for chatter detection, based on PH in combination with supervised learning. Based on the previous work [76], the resulting persistence diagrams are vectorized. The vectorized representation data is then used to train a logistic regression classifier. They report high accuracy in chatter detection.
In a further work (conference [78] and a journal publication [79]), the authors dig deeper into different supervised methods for chatter detection based on topological feature vectors obtained using PH. In these works, different methods on the vectorization of the persistence diagrams are compared and discussed.
Furthermore, in a follow-up work [41], they additionally extend the previous works, by pretraining the methods on a different source domain (turning) and transfer the learned knowledge using transfer learning to a task in the target domain (milling). This work also employs dynamic time warping, to align the time series. Their show-case’s results prove that transfer learning can be used to improve the performance of the chatter, when pretrained on a different dataset.
The task of optimizing the production process of wafers is further addressed in [81]. In the previously discussed work (see Section 5.1), the authors analyze the product itself. In [81], they apply TDA to cluster data about produced wafers to analyze the root cause for defects in the produced goods by only observing latent data from the production process.
A number of variables are measured and recorded for each produced entity, like sensor parameters or electrical measurements. The corresponding high-dimensional data is then analyzed using a proprietary toolset, which identifies homological groups during its processing procedure [^5] The method is described on a very high-level. However, similar to [64], also this work is accounted for PH due to the utilization of homological groups..
A challenge faced in Additive Manufacturing (AM) is to avoid defects in the manufactured parts. Defects can be caused by a variety of reasons, such as the porosity of the material. The work [80] compares three operational strategies for an early stage detection and prediction of part failures due to porosity. In this work, PH is used to analyze process parameters (time series data). The resulting persistence diagram is transformed into PIs, which is then used to train a ML classifier to predict faulty products. Their work concludes by noting that the usage of topological features over basic features from time series data decreases performance in all experimental evaluations. However, this work does not report any experimental evaluation on the combination of both, topological and basic features. Concluding, the authors discuss that the data is incomplete, as these lack some parameters, and the experiments are hard to replicate, due to varying ambient conditions. These results therefore need to be interpreted with caution.
5.3 Manufacturing Engineering
This section covers the applications of TDA in the field of manufacturing engineering. With manufacturing engineering, we refer to the engineering discipline that designs, analyzes, and improves manufacturing processes and systems. The tasks in this area are not centered on the products per se, but rather they are concerned with the overall processes and systems that are used to manufacture these products. These tasks are included within the full manufacturing engineering process (cf. Figure 1) and beyond. Tasks of manufacturing engineering include the optimization of the material flow, the optimization of the production process, and the optimization of the production system, as well as the selection of components and the design of production lines. For this kind of application, $13$ works were found in the literature.
A very common task for a manufacturer is the temporal planning of the production. The demand for a product can vary depending on several factors. Such factors can be the season, the location, the weather, or other events, like promotions or holidays. Not meeting the demand can lead to a loss of customers, while overproduction leads to a monetary loss, e.g., caused by the massive storage or disposal of the products. Depending on the use case, it may be beneficial that similar products are grouped and share forecasting models. In [47], the authors propose a method for demand forecasting with different techniques. In order to generate forecasts for a new product, a predictor-model needs to be selected. For this selection process a $k$-nearest neighbor algorithm based on a Mapper Graph is proposed. By using the topological properties of the historic time-series data, the authors conclude that the selection of the predictor-model is more accurate and much faster than other methods.
The recent work [86] addresses the problem of relying on expert knowledge and experience of machine operators. Turnovers on machinery require re-parameterization of the machines. However, these changes of parameters are based on the experience of the machine operator. This poses the drawback that this re-parameterization is only reproducible up to a certain degree and an operator needs to be trained for a long time in order to gain the required experience on the particular machine type. In their work, the authors propose a strategic synergistic use of existing ML tools to extract a reduced manifold from existing geometric designs of signed distance functions. Using interpolation techniques, the reduced manifolds are used to generate new geometric designs by inferring missing information using clustering techniques. Their work heavily relies on PH and PI.
Material flow optimization is the task of optimizing the schedule and the flow of material through a production system. This task involves transporting raw material from the depot to the production, transporting semi-finished products between the production lines, and transferring finished products from the production to the depot back again. Given the complexity of the production systems, the optimization of the material flow is a challenging task, since all involved components have different capacities, changeover times, and other constraints. Having these constraints in mind, the optimization of the material flow must be covered from a business perspective, as well as from a technical perspective. The task of material flow optimization is a multi-objective optimization problem, where the objectives are to minimize the costs and the time of the material flow. A visual benchmark is proposed in [84]. For this benchmark task, the material flow from depot to the production line is to be optimized as a multi-vehicle routing problem, with the data embodied as a point cloud in a multidimensional space. The evaluation is performed using PH.
An optimization task of another kind is presented in [85]. Given independent moving objects, like grippers or robots within a production environment, these must be secured against collisions with other objects, the surrounding environment, and, most important, human operators. Securing these object can be achieved through physical cages, where the objects operate within. However, these cages can be very complex to built, are inflexible and expensive. A more cost-effective solution is to use virtual cages, where the objects are restricted by a virtual boundary. The task of this work is the syntheses of planar energy-bounded cages that have an optimal configuration for a given object. By means of identifying gripper and force-direction configurations and the application of PH, an optimal configuration is found. For this purpose, the objects and grippers are modeled as point clouds.
Making sure that a product leaving the production line is of a certain quality is a major task in manufacturing engineering. Releasing faulty products can lead to a loss of reputation and, in the worst case, to a loss of human life. To make sure to meet the quality requirements, system-level tests are performed on each product. For the generation of rules to classify the failed parts, [52] proposed a method that employs UMAP for dimensionality reduction. Unlike the applications discussed in Section 5.1, this work does not focus on the detection of discrepancies; rather, it concentrates on generating rules that allow the classification of manufactured parts. The write-up of the paper leaves out details of their proposed solution, but the main ideas seems to be that they embed the information about passed and failed tests in a multidimensional space and then use UMAP to reduce the dimensionality. This representation is then used to generate rules for the classification of failed parts.
OT systems are Cyber-Physical Systems, i.e., where the physical part is controlled by a computer system through sensors and actors. Establishing and extending such systems is a challenging task, as these systems tend to become highly complex and heterogeneous. Identifying repeating patterns allows for the optimization of resource allocation and system configuration, enhancing operational efficiency and reliability. To find repeating patterns in these systems, [88] propose a method to reuse components that are already established in the system, based on labeled graphs. This guarantees higher reliability, lower cost, and lower effort for maintenance. For their method, they use UMAP for dimension reduction.
In Industry 4.0, the security of production systems, and OT systems in general, is a major concern. OT has far different requirements than Information Technology (IT) systems. OT systems, unlike IT systems, directly interact with the physical world, necessitating robust security measures that account for both digital and physical vulnerabilities. This also reflects the challenges in the field of security [97, 33]. Covering the topics on OT security there is only one work found during the survey. In [92], the authors propose using UMAP for downstream tasks in security applications with steel production processes, whereas the data is given as time-series data.
A case study in the herbal medicine manufacturing industry is presented in [89]. In this work, the authors attempt to analyze the degradation of the evaporation process, since this process is a major cost factor in the production. For the analysis, they employ UMAP for dimension reduction on time-series data.
For the task of preventive maintenance, [90] propose a method for the analysis of machine maintenance data. Such datasets are often heterogeneous and multidimensional logs. These data need to be analyzed to find patterns that indicate a failure. In their work, the authors introduce a visual analytics approach for the diagnosis of such heterogeneous and multidimensional machine maintenance data (textual log data), where UMAP is used for dimension reduction as one step in the processing pipeline.
Another approach for the analysis of machine maintenance data is presented in [91]. In their work, the authors mitigate the problem of the degradation of the machinery in the fine blanking industry by observing the wear of the machinery using acoustic emissions. This approach is based on UMAP and hierarchical clustering on time-series data. The authors claim that the data visualized in two dimensions resemble the temporal dependence of the data, while allowing to identify the wear of the machinery.
A use-case on the planning of a multi-robot collective transport tasks is described in [87]. This work uses topological abstractional features for graph reinforcement learning. PH is used to derive higher-order topological information from an input task graph, to facilitate a feature vector for the policy network. The evaluation reports similar performance to state-of-the-art methods, while the necessary computational time was reduced significantly.
The task of monitoring and analyzing the degree of degradation of machine equipment is a crucial step in smart manufacturing. An optimal strategy may prevent machine failures and the overall possible operational time may be maximized. In [93], the authors present a method for the analysis of machine degradation in batch production. In their work, they employ UMAP to cluster time-series data as one step in their analysis pipeline.
The work [94] discusses the tasks of detecting critical events in machinery of mining and metal processing. Critical events may be characterized as patterns within the data, that indicate a potential failure of the machinery. This work proposes the equipment of vibration sensors on the production machinery to record vibrations of the machinery. The data may then be used to improve maintenance strategies, minimize downtimes, or improve production throughput. In their use-case, the acquired data is fourier-transformed and enriched by information from Machine Execution Systems and Enterprise Resource Planning Systems. The resulting set of features is clustered using UMAP.
5.4 Excluded works
Besides being used for data analysis tasks on input data, TDA methods are also used for validating the analysis performed by other methods. For example, in [98] and [99], PH is used for evaluation, and, in [100, 101, 102, 103, 104], and [105], UMAP is used for the same purpose. Since this survey addresses the direct applications of TDA, these works are not included in the results and the remaining discussion.
Furthermore, other articles mention the use of TDA methods on manufacturing systems in their work, but did not do any empirical work. Examples for this are the potential application of TDA methods to hybrid twins for smart manufacturing [106] and 3D printing [29].
6 Discussion
6.1 Interpretation of the Results
As we showed, the number of industrial applications of TDA is rising over the last years, but the number of applications is still low (c.f. Section 5) compared to more established domains, like biology (81) or medicine (84) (numbers according to DONUT [9]). This may result from several factors. First, the discrepancy of focuses: practitioners in industry are primarily engineering-oriented, while TDA experts TDA are more research-focused. Second, the integration of other data-driven methods whose use requires less background, such as ML, into production systems is still in its infancy, and therefore has not yet reached its limits. Third, the lack of mutual awareness. Practitioners in industry are not yet fully informed of the potential of TDA, and TDA experts are not aware of attractiveness of the challenges the industrial and manufacturing setting pose.
Method-wise, we observe a disparity in utilization. The most predominantly utilized method is PH (20 publications). This prevalence might be caused by its interpretability compared to the other methods. The dimension-reduction method UMAP is used in 11 referenced works. All of them underscore UMAPs topology-preserving characteristics, favoring its use over other dimension-reduction methods. The method used the least is Mapper with only 4 works. We speculate that a reason for this is that Mapper has been patented and commercialized [107], possibly preventing the academic community to freely study its applications. At the same time, the need to protect a TDA method proofs its commercial value.
UMAP and PH are the methods that are applied on the most diverse data types. PH is utilized on time-series data, point clouds, and scalar fields. Likewise, also UMAP is used on a variety of data types, including time-series data, point clouds, and even log data. Interestingly, Mapper is currently applied to time-series data only, except one work dealing with scalar fields.
Unlike the others, Section 5.2 mostly (with one exception) leverages time series data. This is rather natural, since on process level is the level where time series data is most prevalent. However, the information about the process itself can be also embedded in a different format, e.g., as summary information of a production process, similar to what is being offered by the OPC UA specification [108] for Injection Moulding Machines to MES systems. In this case, the process information could also be embedded as a graph or a point cloud. This approach can be observed in the only, relatively recent work that is not dealing with time series data.
The application of UMAP is absent in Cluster B, where no studies implementing dimension reduction have been identified – Mapper is instead used for clustering purposes. However, exploring dimension reduction techniques, such as UMAP, could prove beneficial in the context of multivariate time-series data, too.
6.2 Opportunities and Challenges
We now list some pros and cons of the described TDA tools.
Opportunities for Application
UMAP helps against the curse of dimensionality, a prevalent challenge in data analysis characterized by the sparsity of data in high-dimensional spaces.
PH provides novel techniques to automate analytical processes by considering the full value range of a parameter. This approach is highlighted, for example, in the findings of a survey presented in [43].
UMAP and Mapper can deal with large datasets, a particularly valuable property in industrial settings, where data are usually big.
In general, TDA methods are more transparent than, for example, ML, where the analysis happens in a “black box”. Businesses typically prefer methods that are reliable over those that might offer better accuracy but are less understandable. From a strategic standpoint, this makes TDA a compelling choice for the application in industrial settings.
Moreover, TDA, by its very nature, extracts or uses information about shapes and topological features that no other data analysis method can find. However, one does not need to restrict to TDA, and it can be seamlessly integrated with other methodologies. An illustrative example is its combination with techniques fromML, where topological features are extracted using techniques from TDA. According to [21], the application of TDA with ML can be grouped into two main categories, namely intrinsic- and extrinsic topological features for ML. Intrinsic topological features incorporate topological analysis of the machine learning model itself and influence the model’s architecture or training. Interestingly, none of the identified works of this survey is of this category. We see this gap as potential opportunity for future TDA research in this application domain. Extrinsic topological features, however, enable the use of topological features extracted from a given dataset. This is enabled by vectorized persistence diagrams, like PI [54] or specialized kernels [56]. The degree of facilitating the topological information can vary and highly depends on the use-case and data:
  • 1.
    Topological features are sufficient on its own, i.e., naive and direct analysis of vectorized representations of persistence diagrams. An example from this survey is the supervised analysis of chatter on time-series signals [78].
  • 2.
    Adding topological information to existing pipelines. Information about shape is often informative and topological features can be used as complementary features to improve existing methods. The work in [87] illustrates this on a use-case by adding topological features to an already existing method and reporting an improved performance.
Another noteworthy approach to combine ML and TDA is, that ML methods preprocess data, so that topological aspects can be derived from the resulting representation. As an example, [72] facilitate a vision transformer to extract higher-level features from wafer maps, which are then analyzed using Mapper.
Lastly, TDA software are usually open source and thus very accessible. With versions available in programming languages such as Python, R, and C++, TDA methods can be applied almost off-the-shelf, without a deep understanding of the theoretical underpinnings.
Open Challenges in TDA
Nonetheless, there are many open challenges in TDA. For example, despite its success, and unlike Mapper and UMAP, PH does not scale very well to large datasets. This is particularly challenging since the method itself is hard to parallelize.
Moreover, many applications have several parameters that need to be considered simultaneously, and the so-called “multiparameter persistence” is starting only now to have viable software [109].
In addition to these limitations, there are also some challenges in the application of TDA in the industrial setting.
Challenges in Application
Although implementations of Mapper are freely available, its application faces a significant obstacle due to patent protection [107]. This legal constraint could hinder its adoption in industrial and commercial settings, unlike other methodologies that do not encounter legal barriers.
Furthermore, while methods such as ML can be applied in a problem-agnostic manner, the use of TDA methods require a more discerning approach. A deep understanding of the data and the specific problem at hand is crucial for employing topological methods effectively, making them less straightforward.
Another difficulty is the selection of the appropriate TDA tool for the task at hand. The descriptions and discussion in Section 3 and Section 5 help mitigate this issue. To further address it, we provide some guidelines tailored to industry experts to select the appropriate TDA method based on their specific data type or analysis requirements. These guidelines also offer an overview of some possible future research directions. We are not aware of a similar list of guideline.
  1. 1.
    Time Series Data: For analyzes involving time series, PH is recommended due to its effectiveness in capturing temporal topological features. For dimension-reduction of multivariate time-series data, UMAP is advisable, as it preserves the topological structure of the data.
  2. 2.
    Point Cloud Data: If the point cloud is in low ambient dimension (up to 2) then PH, with the so-called alpha complex, is extremely efficient also for big datasets ($10^{4}$ points). In higher ambient dimension, PH is only feasible, so far, for datasets of $10^{2}$ or $10^{3}$ points, especially if one is interested in topological features of higher dimension (such as cavities), and not only low dimension (such as connected components and loops). For larger datasets, Mapper or UMAP should be considered first. The choice between them should be guided by the specific characteristics and requirements of the data.
  3. 3.
    Automation of Processes: For the automation of tasks or processes, anytime there is a clear choice for a varying parameter PH should be considered.
  4. 4.
    High-Dimensional Data: For data residing in high ambient dimensions, UMAP is the preferred method to mitigate the curse of dimensionality.
This list is an introductory, practical guide for industry professionals. It is incomplete and possibly suboptimal once one is more familiar with TDA tools, but it gives the beginner very useful rule-of-thumb to approach the field.
We hope, moreover, that it will encourage practitioners in areas where PH, UMAP, or Mapper have yet to be used to test these methods. We anticipate that several tasks can be analyzed with TDA methods, resulting novel insights in these areas. Even problems considered “solved” (refer to [110] for examples and datasets) can be re-examined and further explored.
6.3 Future Research Directions
The realm of Industry 4.0 describes the utilization of advanced industrial analytics for adaptive decision-making. Despite this focus, this survey has not identified any studies on complex decision-making tasks such as automated control or data-driven decision-making. However, achieving the objective of adaptive manufacturing systems requires sophisticated methods. We claim that TDA can be effectively utilized for this purpose, thereby facilitating the realization of smart manufacturing systems.
In the setting of industrial production and manufacturing, the scale and physical nature of data are crucial aspects. This contrasts with fields like finance, where data is predominantly abstract, encompassing diverse physical measurements such as pressures, temperatures, and positions. Consequently, transforming features in this industry can significantly affect the data’s topological and geometrical characteristics. Nevertheless, employing methods that preserve topology, such as UMAP, to reduce data dimensionality can maintain the integrity of the data’s topological and geometrical structure. This strategy is particularly promising for handling multi-variate time series data common in industrial production processes.
TDA can also be applied on data that has been transformed by the Fourier Transform. Analyzing data in its Fourier-transformed representation is a common practice across various industrial sectors outside TDA, including control theory and image processing. To the author’s knowledge, there is limited literature on this specific application within the TDA field. The only identified publication is by Huber, which presents a toy example focusing on the feedback optimization of a closed-loop controller [39].
As discussed in Section 6.1, the application of TDA in industrial production and manufacturing is still in its early stages. Thus, there is a substantial opportunity for cross-disciplinary collaboration between the fields of production, manufacturing, and TDA. Engaging in such collaborations is expected to generate novel applications and methodologies, providing benefits to both areas.
Lastly, we would like to see, with a few years of distance, a recurring survey on the same topic as this work. This would allow to observe the development of the application of TDA in this domain and give insights into the development of the field. Given the nature of the method used in this work, this is highly reproducible and open for other researchers to extend this work.
6.4 Limitations of this Study
The results of this study are biased towards the chosen method. First, the application of TDA in smart manufacturing is a growing field. Thus, this survey can only be seen as a snapshot of the current published scientific literature. The search terms for this study are intentionally broad, followed by a very restrictive manual filtering procedure. However, these search terms may not capture the full literature available. Moreover, since we restricted to some publishers and databases, some publications may have been missed, although a semi-exhaustive preliminary search was conducted to mitigate this bias. Furthermore, gray literature, like pre-prints, is also ignored for this study.
We also want to highlight that this study only considers works in industrial manufacturing. There are other industrial fields that may also interest this article’s target audience. There are, for example, works on process optimization in oil, gas, and chemical industry that employ methods from TDA (see [43, 111, 112]).
Lastly, we want to emphasize that not all industrial applications are published in academic literature. Revealing the techniques employed on proprietary issues frequently results in a competitive edge. In the authors’ experience, these applications are sometimes published only after a significant delay, if at all.
7 Conclusion
This study contributes to the existing body of knowledge in several aspects:
  1. 1.
    We provide an overview of the state of the art in the application of TDA within industrial and manufacturing systems;
  2. 2.
    Based on the literature, we outline the areas where TDA methods have been implemented successfully, showcasing the diversity and potential of TDA in addressing complex industrial challenges;
  3. 3.
    The identified application scenarios are thoroughly analyzed, revealing existing gaps and potential future research directions, not confined to the presented scenarios but applicable in a broader context;
  4. 4.
    Finally, we provide a set of guidelines to aid practitioners and researchers in applying TDA within the field of industrial production and manufacturing.
For the survey, a transparent and rigorous methodology was employed to search for and identify relevant literature, ensuring the reproducibility of our findings. This approach yielded 35 works, which were analyzed. These studies were categorized into three distinct groups: Quality Control at the Product Level, Quality Control at the Process Level, and Manufacturing Engineering. Each study is reviewed, focusing on the data formats used and the specific TDA methods applied to each case.
Through this research, we demonstrate that TDA is exceptionally well-suited for analyzing complex datasets derived from sensors and other devices within the realm of industrial production and manufacturing. Additionally, we highlight that the application of TDA in this domain is still nascent, presenting significant potential for future research. To address the challenges of integrating TDA into industrial production, we propose a set of guidelines to aid practitioners in applying TDA within this field.
The evolution of TDA within the context of industrial production and manufacturing is at a preliminary stage. Despite existing challenges, we are optimistic about the substantial potential of TDA in this domain. This work serves as an initial step towards raising awareness of this potential. Looking ahead, we envision a robust collaboration between academia and industry to nurture the field’s growth and expand the range of applications.
Acknowledgements
Martin Uray and Stefan Huber are supported by the Christian Doppler Research Association (JRC ISIA). Barbara Giunti and Michael Kerber were partially or fully supported by the Austrian Science Fund (FWF) [P 33765-N].
Author contributions
Martin Uray: Conceptualization, Methodology, Software, Validation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Project administration. Barbara Giunti: Writing - Original Draft, Writing - Review & Editing, Visualization. Michael Kerber: Writing - Original Draft, Writing - Review & Editing. Stefan Huber: Writing - Original Draft, Writing - Review & Editing, Funding acquisition, Supervision.